dovbushbeatz:

Help please Sydney made the largest s'more when camping with 70g of marshmallows, 50g of graham cracker and 100g of chocolate. To find the percent composition of the chocolate, which equation would she use? Question 1 options: 100g/120g*100 100g/220g*100 220g/100g*100 120g/100g*100

jhonyy9:

70g+50g+100g = 220g using the rule of 3 simple you can write 220g mean 100% 100g mean x cross multiplie and calcule x value x = 100*100/220 = ? %

jhonyy9:

@Laylalyssa